In a statistical study relating to the prices (in T) of two shares, X and Y, the following two regression lines were found as 8X - 10Y + 70 = 0 and 20X - 9Y - 65 = 0. The standard deviation of X = 3, then find (i) the values of X and Y, (ii) r(X, Y), and (iii) standard deviation of Y.


1
Expert's answer
2022-06-14T12:07:48-0400

(i) The given equation of the lines of regression are


"8X - 10Y + 70 = 0""20X - 9Y - 65 = 0"


"40X - 50Y + 350 = 0""40X - 18Y - 130 = 0"

"-32Y+480=0"

"Y=15"

"8X - 10(15) + 70 = 0"

"X=10"

"X=10, Y=15"


(ii) Let the equation "8X - 10Y + 70 = 0" be the regression equation of "Y" on "X." Then


"Y=0.8X+7"

Comparing it with "Y=b_{YX}X+a," we get "b_{YX}=0.8."

Let the equation "20X - 9Y - 65 = 0" be the regression equation of "X" on "Y." Then


"X=0.45X+3.25"

Comparing it with "X=b_{XY}Y+a'," we get "b_{XY}=0.45."


"r=\pm\sqrt{b_{XY}\cdot b_{YX}}"

Since "b_{XY},b_{YX}" both are positive, then "r" is also positive:


"r=\sqrt{0.8\cdot 0.45}=0.6"

(iii)


"b_{YX}=r\dfrac{\sigma_Y}{\sigma_X}"

Then


"\sigma_Y=b_{YX}\dfrac{\sigma_X}{r}"

"\sigma_Y=0.8(\dfrac{3}{0.6})=4"